The demand function of a firm is 


𝑄=90−1,5𝑃,

Q=90−1,5P,

where 𝑃

P and 𝑄

Q represent the price and quantity, respectively. At what price is revenue a maximum?

a.15

b.30

c.270

d.90


1
Expert's answer
2022-03-25T13:28:55-0400

The revenue is at maximum when TR'(Q) = MR = 0, so:

P = 60 - 2/3Q,

MR = 60 - 4/3Q = 0,

Q = 45 units,

P = 60 - 30 = 30.

So, the correct answer is b.